# Mixed precision iterative refinement for least squares with linear equality constraints and generalized least squares problems

### v2

@@ 4.2 Analysis for the GRQ/GQR factorization
Let $A\in\mathbb{C}^{m\times n}$ and $B\in\mathbb{C}^{p\times n}$. Suppose $(\hat{R},\hat{T})$ are the computed triangular matrices from the GRQ factorization of $(B,A)$ via Householder transformations. Then there exist matrices $\Delta E_{1}$, $\Delta E_{2}$, and unitary matrices $Q$, $Z$ such that
@@
$\big{\lVert}\operatorname{f{}l}(AQ^{*})-AQ^{*}\big{\rVert}_{\mathsf{F}}\leq\epsilon_{\mathtt{unmqr}}(n)\lVert A\rVert_{\mathsf{F}},$ (4.1)
@@
$\big{\lVert}Z\hat{T}-\operatorname{f{}l}(AQ^{*})\big{\rVert}_{\mathsf{F}}\leq\epsilon_{\mathtt{qr}}(m,n)\lVert\operatorname{f{}l}(AQ^{*})\rVert_{\mathsf{F}}\leq\epsilon_{\mathtt{qr}}(m,n)(1+\epsilon_{\mathtt{unmqr}}(n))\lVert A\rVert_{\mathsf{F}},$ (4.2)
@@
Let $W\in\mathbb{C}^{n\times m}$ and $V\in\mathbb{C}^{n\times p}$. Suppose $(\hat{R},\hat{T})$ are the computed triangular matrices from the GQR factorization of $(W,V)$ via Householder transformations. Then there exist matrices $\Delta F_{1}$, $\Delta F_{2}$, and unitary matrices $Q$, $Z$ such that
$W+\Delta F_{1}=Q\begin{bmatrix}R\\ 0\end{bmatrix},\qquad V+\Delta F_{2}=QTZ,$
@@
$\lVert\Delta F_{2}\rVert_{\mathsf{F}}\leq\big{(}\epsilon_{\mathtt{unmqr}}(n)+\epsilon_{\mathtt{qr}}(p,n)+\epsilon_{\mathtt{unmqr}}(n)\epsilon_{\mathtt{qr}}(p,n)\big{)}\lVert V\rVert_{\mathsf{F}}.$
@@ 4.3 Analysis for the LSE problem
and the following bounds are satisfied.
@@
where $\eta_{0}$, $\eta_{1}$, and $\eta_{2}$ are small constants of $m$, $n$, $p$, and the unit roundoffs:
@@
$\eta_{2}$ $=\max\bigl{\{}\epsilon_{\mathtt{trsv}}^{s}(n-p),\epsilon_{\mathtt{gemv}}^{s}(n-p),\epsilon_{\mathtt{gemv}}^{s}(m-n+p)\bigr{\}}.$
@@
$\tilde{F}=\begin{bmatrix}\alpha I_{m}&0&A\\ 0&0&\beta B\\ A^{*}&\beta B^{*}&0\end{bmatrix},\quad\tilde{u}=\begin{bmatrix}r\\ -v\\ x\end{bmatrix},\quad\Delta\tilde{u}=\begin{bmatrix}\alpha^{-1/2}\Delta r\\ -\alpha^{-1/2}\beta^{-1}\Delta v\\ \alpha^{1/2}\Delta x\end{bmatrix},\quad\tilde{s}=\begin{bmatrix}\alpha^{1/2}f_{1}\\ \alpha^{1/2}\beta f_{2}\\ \alpha^{-1/2}f_{3}\end{bmatrix},$
where $\alpha$ and $\beta$ are scaling parameters. Then system (3.2) is equivalent to $\tilde{F}\Delta\tilde{u}=\tilde{s}$, and the refinement stage of Algorithm 2 can be regarded as a solver of the linear system $\tilde{F}\Delta\tilde{u}=\tilde{s}$. We have the following result.
@@
The proof is similar to [11, Section 2] and thus we omit it. ∎
@@ 4.4 Analysis for the GLS problem
and the following bounds are satisfied.
@@
where $\eta_{0}$, $\eta_{1}$, and $\eta_{2}$ are small constants of $m$, $n$, $p$, and the unit roundoffs:
@@
$\eta_{2}$ $=\max\big{\{}\epsilon_{\mathtt{trsv}}^{s}(n-m),\epsilon_{\mathtt{gemv}}^{s}(n-m),\epsilon_{\mathtt{gemv}}^{s}(p-n+m)\big{\}}.$
@@
$\tilde{F}=\begin{bmatrix}I_{p}&V^{*}&0\\ V&0&W\\ 0&W^{*}&0\end{bmatrix}\qquad\text{and}\qquad\tilde{u}=\begin{bmatrix}y\\ -z\\ x\end{bmatrix}.$
For the GLS problem, the iterative refinement method is also independent of scaling. To keep things straightforward, we do not cover the scaling discussion here again. Similarly to Theorem 4.8, the limiting accuracy of the GLS problem is presented as follows.
@@
Theorems 4.8 and 4.10 imply that $\kappa_{\infty}(\tilde{F})$ must be bounded by $\mathcal{O}(\bm{u}_{f}^{-1})$ in order for the algorithm to converge, and precision $\bm{u}$ together with the product $\operatorname{cond}(\tilde{F},\tilde{u})\bm{u}_{r}$ determine the attainable accuracy. A specific example of that can be found in Table 2 as in [10] by employing IEEE half/single/double precisions.
@@
$\bm{u}_{f}$ $\bm{u}$ $\bm{u}_{r}$ Upper bound on $\kappa_{\infty}(\tilde{F})$ Forward error
@@
Another conclusion of Theorems 4.8 and 4.10 is that, if we use a higher (than working) precision to compute the residual in the iterative refinement, i.e., $\bm{u}_{r}\ll\bm{u}$, then $\operatorname{cond}(\tilde{F},\tilde{u})\bm{u}_{r}$ may no longer be the leading error term in (4.3) and (4.4). The forward error is thus reduced from $\operatorname{cond}(\tilde{F},\tilde{u})\cdot 10^{-16}$ to double precision; as illustrated by certain scenarios in Table 2. In other words, the mixed precision algorithm improves the accuracy of the solution.
@@ 5 GMRES-based iterative refinement
As shown in Section 4, the mixed precision classical iterative refinement algorithm has the obvious drawback that it only works for the case $\bm{u}_{f}\kappa_{\infty}(\tilde{F})\lesssim 1$. To overcome this problem, we use preconditioned GMRES algorithms to solve the saddle-point system. Since the GRQ or GQR factorization is already obtained, we may construct the preconditioners using these computed factors. For a brief introduction to the GMRES theory, see [33].
Let $\tilde{F}$ be the augmented matrix with a scaling parameter $\alpha$, i.e.,
[…]
As discussed in Remark 4.2, for GMRES-based iterative refinement, we shall use $\beta$ to keep the norms of $A$ and $B$ (or $V$ and $W$) at the same magnitude before the GRQ or GQR factorization, and meanwhile employ $\alpha$ to reduce the condition number of $\tilde{F}$.
These two augmented systems with $\tilde{F}$ can be regarded as symmetric saddle point systems. In [4, 32], both left and block diagonal preconditioners for saddle point systems are studied. Here we propose an improved version of preconditioners for addressing LSE and GLS problems.
@@ 5.1 Left preconditioners
A natural choice of the left preconditioner for the LSE problem, inspired by directly calculating the pseudoinverse of $\tilde{F}$, is given as
$M=\begin{bmatrix}\alpha^{-1}I_{m}&-\alpha^{-1}AB^{\dagger}&0\\ -\alpha^{-1}(AB^{\dagger})^{*}&\alpha^{-1}(B^{\dagger})^{*}A^{*}AB^{\dagger}&(B^{\dagger})^{*}\\ 0&B^{\dagger}&0\end{bmatrix},$ (5.1)
where $B^{\dagger}$ is the Moore–Penrose pseudoinverse of $B$.
@@
$M=\begin{bmatrix}\alpha^{-1}I_{p}&0&-\alpha^{-1}V^{*}(W^{\dagger})^{*}\\ 0&0&(W^{\dagger})^{*}\\ -\alpha^{-1}W^{\dagger}V&W^{\dagger}&\alpha^{-1}W^{\dagger}VV^{*}(W^{\dagger})^{*}\end{bmatrix}.$ (5.2)
For both LSE and GLS problems, applying the preconditioner $M$ to a vector is done using the procedure in Section 3 in practice.
The convergence of the GMRES method with these left preconditioners follows from the arguments in [11]. However, as shown in Section 6, these specific left preconditioners do not work well in the mixed precision GMRES-based iterative refinement algorithm and sometimes take an awfully long time to converge as each iteration is numerically very expensive. This motivates us to develop other preconditioners in order to speed up the process.
@@ 5.2 Block-diagonal split preconditioner for the LSE problem
$\begin{split}M_{l}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}S\hat{R}^{-1},\alpha^{1/2}\hat{T}_{1}^{-*}Q\big{\}},\\ M_{r}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}\hat{R}^{-*}S^{*},\alpha^{1/2}Q^{*}\hat{T}_{1}^{-1}\big{\}},\end{split}$ (5.3)
@@
$X=\begin{bmatrix}I_{m}&0&Z_{1}\\ 0&0&\bar{I}_{p}\\ Z_{1}^{*}&\bar{I}_{p}^{*}&0\end{bmatrix},\qquad\Delta E=M_{l}\begin{bmatrix}0&0&\Delta E_{2}\\ 0&0&\Delta E_{1}\\ \Delta E_{2}^{*}&\Delta E_{1}^{*}&0\end{bmatrix}M_{r},$
@@
$M_{l}\tilde{F}M_{r}$ $=M_{l}\begin{bmatrix}\alpha I_{m}&0&Z\hat{T}Q\\ 0&0&[0,\hat{R}]Q\\ (Z\hat{T}Q)^{*}&\big{(}[0,\hat{R}]Q\big{)}^{*}&0\end{bmatrix}M_{r}-M_{l}\begin{bmatrix}0&0&\Delta E_{2}\\ 0&0&\Delta E_{1}\\ \Delta E_{2}^{*}&\Delta E_{1}^{*}&0\end{bmatrix}M_{r}=X-\Delta E,$
where $\lVert\Delta E\rVert_{2}\leq\sqrt{2}\,\lVert M_{l}\rVert_{2}\bigl{(}\lVert\Delta E_{1}\rVert_{\mathsf{F}}+\lVert\Delta E_{2}\rVert_{\mathsf{F}}\bigr{)}\lVert M_{r}\rVert_{2}.$ By Lemma 4.5, $\lVert\Delta E_{1}\rVert_{\mathsf{F}}=\mathcal{O}(\bm{u}_{f})\lVert B\rVert_{2}$ and $\lVert\Delta E_{2}\rVert_{\mathsf{F}}=\mathcal{O}(\bm{u}_{f})\lVert A\rVert_{2}$. Note that from (5.3), we have
$\lVert M_{l}\rVert_{2}=\lVert M_{r}\rVert_{2}\leq\alpha^{-1/2}+\bigl{(}\alpha^{-1/2}+\mathcal{O}(\bm{u}_{f})\bigr{)}\lVert A\rVert_{2}\lVert B^{\dagger}\rVert_{2}+\bigl{(}\alpha^{1/2}+\mathcal{O}(\bm{u}_{f})\bigr{)}\lVert A^{\dagger}\rVert_{2}.$
Therefore
$\lVert\Delta E\rVert_{2}\leq\big{(}\alpha^{-1/2}+\alpha^{-1/2}\lVert A\rVert_{2}\lVert B^{\dagger}\rVert_{2}+\alpha^{1/2}\lVert A^{\dagger}\rVert_{2}\big{)}^{2}\bigl{(}\lVert A\rVert_{2}+\lVert B\rVert_{2}\bigr{)}\mathcal{O}(\bm{u}_{f})\equiv\Phi.$
The next step is to derive the eigenvalues of $X$. A similar analysis can be found in [4, Sect. 10]. By examining the specific structure of the preconditioned matrix, we verify that $X$ satisfies
@@
A detailed proof of this property is presented in Lemma C.1. By [16, Corollary 2.4.4], as long as $2\Phi<\lambda_{2}$, we have
@@
By the norm inequalities [20, Table 6.2],
@@
$\begin{split}M_{l}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}Y\hat{R}^{-1},\alpha^{1/2}U^{-*}Q\big{\}},\\ M_{r}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}\hat{R}^{-*}Y^{*},\alpha^{1/2}Q^{*}U^{-1}\big{\}},\end{split}$ (5.4)
@@
where $\bar{Z}=[Z,0]\in\mathbb{C}^{m\times n}$. Similarly, one can prove that the condition number of the preconditioned matrix satisfies
@@
Although it does not necessarily ensure a fast convergence rate theoretically, the condition number of $M_{l}\tilde{F}M_{r}$ has strong implications on the magnitude of the backward and forward errors for the GMRES-based algorithms. For references, a comprehensive bound for the backward and forward errors in the FGMRES algorithm is provided in [8, Sect. 2]. In practice, the block-diagonal split two-sided preconditioner, due to its smaller condition number and reduced communication cost, shows superior performance in numerical experiments, and therefore emerges as a popular choice for GMRES-based iterative refinement.

### v3

@@ 4.2 Analysis for the GRQ/GQR factorization
[…]
@@
$\big{\lVert}\operatorname{f{}l}(AQ^{*})-AQ^{*}\big{\rVert}_{\mathsf{F}}\leq\epsilon_{\mathtt{unmqr}}(n)\lVert A\rVert_{\mathsf{F}}.$ (4.1)
@@
$\big{\lVert}Z\hat{T}-\operatorname{f{}l}(AQ^{*})\big{\rVert}_{\mathsf{F}}\leq\epsilon_{\mathtt{qr}}(m,n)\lVert\operatorname{f{}l}(AQ^{*})\rVert_{\mathsf{F}}\leq\epsilon_{\mathtt{qr}}(m,n)(1+\epsilon_{\mathtt{unmqr}}(n))\lVert A\rVert_{\mathsf{F}}.$ (4.2)
@@
[…]
$W+\Delta F_{1}=Q\begin{bmatrix}\hat{R}\\ 0\end{bmatrix},\qquad V+\Delta F_{2}=Q\hat{T}Z,$
@@
$\lVert\Delta F_{2}\rVert_{\mathsf{F}}\leq\bigl{(}\epsilon_{\mathtt{unmqr}}(n)+\epsilon_{\mathtt{qr}}(p,n)+\epsilon_{\mathtt{unmqr}}(n)\epsilon_{\mathtt{qr}}(p,n)\bigr{)}\lVert V\rVert_{\mathsf{F}}.$
@@ 4.3 Analysis for the LSE problem
and the following bounds are satisfied:
@@
where
@@
$\eta_{2}$ $=\max\bigl{\{}\epsilon_{\mathtt{trsv}}^{s}(n-p),\epsilon_{\mathtt{gemv}}^{s}(n-p),\epsilon_{\mathtt{gemv}}^{s}(m-n+p)\bigr{\}}$
are small constants of $m$, $n$, $p$, and the unit roundoffs.
@@
$\tilde{F}=\begin{bmatrix}\alpha I_{m}&0&A\\ 0&0&\beta B\\ A^{*}&\beta B^{*}&0\end{bmatrix},\quad\tilde{u}=\begin{bmatrix}r\\ -v\\ x\end{bmatrix},\quad\Delta\tilde{u}=\begin{bmatrix}\alpha^{-1/2}\Delta r\\ -\alpha^{-1/2}\beta^{-1}\Delta v\\ \alpha^{1/2}\Delta x\end{bmatrix},\quad\tilde{s}=\begin{bmatrix}\alpha^{1/2}f_{1}\\ \alpha^{1/2}\beta f_{2}\\ \alpha^{-1/2}f_{3}\end{bmatrix}$, (4.3)
where $\alpha$ and $\beta$ are positive scaling parameters. Then system (3.2) is equivalent to $\tilde{F}\Delta\tilde{u}=\tilde{s}$, and the refinement stage of Algorithm 2 can be regarded as a solver of the linear system $\tilde{F}\Delta\tilde{u}=\tilde{s}$. We have the following result.
@@
The proof is similar to [13, Section 2] and thus we omit it. ∎
@@ 4.4 Analysis for the GLS problem
and the following bounds are satisfied:
@@
where
@@
$\eta_{2}$ $=\max\big{\{}\epsilon_{\mathtt{trsv}}^{s}(n-m),\epsilon_{\mathtt{gemv}}^{s}(n-m),\epsilon_{\mathtt{gemv}}^{s}(p-n+m)\big{\}}$
are small constants of $m$, $n$, $p$, and the unit roundoffs.
@@
$\tilde{F}=\begin{bmatrix}\alpha I_{p}&V^{*}&0\\ V&0&\beta W\\ 0&\beta W^{*}&0\end{bmatrix},\quad\tilde{u}=\begin{bmatrix}y\\ -z\\ x\end{bmatrix},\quad\Delta\tilde{u}=\begin{bmatrix}\alpha^{-1/2}\Delta y\\ -\alpha^{1/2}\Delta z\\ \alpha^{-1/2}\beta^{-1}\Delta x\end{bmatrix},\quad\tilde{s}=\begin{bmatrix}\alpha^{1/2}f_{1}\\ \alpha^{-1/2}f_{2}\\ \alpha^{1/2}\beta f_{3}\end{bmatrix}$, (4.5)
where $\alpha$ and $\beta$ are positive scaling parameters. Then $\tilde{F}\Delta\tilde{u}=\tilde{s}$. Similarly to Theorem 4.8, the limiting accuracy of the GLS problem is presented as follows.
@@
[…]
@@
[…]
@@
Another conclusion of Theorems 4.8 and 4.10 is that, if we use a higher (than working) precision to compute the residual in the iterative refinement, i.e., $\bm{u}_{r}\ll\bm{u}$, then $\operatorname{cond}(\tilde{F},\tilde{u})\bm{u}_{r}$ may no longer be the leading error term in (4.4) and (4.6). The forward error is thus reduced from $\operatorname{cond}(\tilde{F},\tilde{u})\cdot 10^{-16}$ to double precision, as illustrated by certain scenarios in Table 2. In other words, the mixed precision algorithm improves the accuracy of the solution.
@@ 5 GMRES-based iterative refinement
As shown in Section 4, the mixed precision classical iterative refinement algorithm has the obvious drawback that it only works for the case $\bm{u}_{f}\kappa_{\infty}(\tilde{F})\lesssim 1$. To overcome this problem, we use preconditioned GMRES algorithms to solve the saddle-point system. Since the GRQ or GQR factorization is already obtained, we may construct the preconditioners using these computed factors. For a brief introduction to the GMRES algorithm, see [35].
Let $\tilde{F}$ be the augmented matrix with scaling parameters $\alpha$ and $\beta$, i.e.,
$\tilde{F}=\begin{bmatrix}\alpha I_{m}&0&A\\ 0&0&\beta B\\ A^{*}&\beta B^{*}&0\end{bmatrix}\quad\text{(LSE)}\qquad\text{or}\qquad\tilde{F}=\begin{bmatrix}\alpha I_{p}&V^{*}&0\\ V&0&\beta W\\ 0&\beta W^{*}&0\end{bmatrix}\quad\text{(GLS)}.$
Here the parameters $\alpha$ and $\beta$ are designed to reduce the condition number of $\tilde{F}$ and to balance the norms of matrices $A$ and $B$ for the LSE problem ($V$ and $W$ for the GLS problem). Determining the optimal values of $\alpha$ and $\beta$ for the LSE and GLS problems is theoretically challenging, as the spectrum of $\tilde{F}$ lacks clarity. For LSE problems, during numerical experiments, we set $\beta\approx\lVert A\rVert_{2}/\lVert B\rVert_{2}$ to achieve norm balance, and our choice of $\alpha$ is derived from the particular scenario $B=0$, where the LSE problem simplifies to the standard least squares problem. Under the assumption $B=0$, the optimal choice of $\alpha$ is $\alpha=2^{-1/2}\sigma_{\min}(A)$ [6, 10]. Since computing $\sigma_{\min}(A)$ is costly, we instead use $\alpha=\lVert r_{0}\rVert_{2}$, a choice inspired by [4] in our numerical experiments. Similarly, we set $\alpha=\lVert y_{0}\rVert_{2}$ and $\beta\approx\lVert V\rVert_{2}/\lVert W\rVert_{2}$ for GLS problems.
These two augmented systems with $\tilde{F}$ can be regarded as symmetric saddle-point systems. In [5, 34], both left and block diagonal preconditioners for saddle-point systems are studied. Here we propose an improved version of preconditioners for addressing LSE and GLS problems.
@@ 5.1 Left preconditioners
A natural choice of the left preconditioner for the LSE problem, inspired by directly calculating the inverse of $\tilde{F}$, is given as
$M=\begin{bmatrix}\alpha^{-1}I_{m}&-\alpha^{-1}\beta^{-1}AB^{\dagger}&0\\ -\alpha^{-1}\beta^{-1}(AB^{\dagger})^{*}&\alpha^{-1}\beta^{-2}(B^{\dagger})^{*}A^{*}AB^{\dagger}&\beta^{-1}(B^{\dagger})^{*}\\ 0&\beta^{-1}B^{\dagger}&0\end{bmatrix}$. (5.1)
@@
$M=\begin{bmatrix}\alpha^{-1}I_{p}&0&-\alpha^{-1}\beta^{-1}V^{*}(W^{\dagger})^{*}\\ 0&0&\beta^{-1}(W^{\dagger})^{*}\\ -\alpha^{-1}\beta^{-1}W^{\dagger}V&\beta^{-1}W^{\dagger}&\alpha^{-1}\beta^{-2}W^{\dagger}VV^{*}(W^{\dagger})^{*}\end{bmatrix}$. (5.2)
For both LSE and GLS problems, applying the preconditioner $M$ to a vector is conducted using the procedure in Section 3 in practice.
The convergence of the GMRES algorithm with these left preconditioners follows from the arguments in [13]. However, as shown in Section 6, these specific left preconditioners do not work well in the mixed precision GMRES-based iterative refinement algorithm and sometimes take an awfully long time to converge as each iteration is numerically very expensive. This motivates us to develop other preconditioners in order to speed up the process.
@@ 5.2 Block-diagonal split preconditioner for the LSE problem
$\begin{split}M_{l}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}\beta^{-1}S\hat{R}^{-1},\alpha^{1/2}\hat{T}_{1}^{-*}Q\big{\}},\\ M_{r}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}\beta^{-1}\hat{R}^{-*}S^{*},\alpha^{1/2}Q^{*}\hat{T}_{1}^{-1}\big{\}},\end{split}$ (5.3)
@@
$X=\begin{bmatrix}I_{m}&0&Z_{1}\\ 0&0&\bar{I}_{p}\\ Z_{1}^{*}&\bar{I}_{p}^{*}&0\end{bmatrix},\qquad\Delta E=M_{l}\begin{bmatrix}0&0&\Delta E_{2}\\ 0&0&\beta\Delta E_{1}\\ \Delta E_{2}^{*}&\beta\Delta E_{1}^{*}&0\end{bmatrix}M_{r},$
@@
$M_{l}\tilde{F}M_{r}=X-\Delta E.$
By Lemma 4.5, $\lVert\Delta E_{1}\rVert_{2}=\mathcal{O}(\bm{u}_{f})\lVert B\rVert_{2}$ and $\lVert\Delta E_{2}\rVert_{2}=\mathcal{O}(\bm{u}_{f})\lVert A\rVert_{2}$. Note that from (5.3), we have
$\lVert\Delta E\rVert_{2}=\left\lVert\begin{bmatrix}0&0&\Delta E_{2}Q^{*}\hat{T}_{1}^{-1}\\ 0&0&S\hat{R}^{-1}\Delta E_{1}Q^{*}\hat{T}_{1}^{-1}\\ \hat{T}_{1}^{-*}Q\Delta E_{2}^{*}&\hat{T}_{1}^{-*}Q\Delta E_{1}^{*}\hat{R}^{-*}S^{*}&0\end{bmatrix}\right\rVert_{2}\leq\mathcal{O}(\bm{u}_{f})\kappa_{2}(A)\kappa_{2}(B).$
The next step is then to derive the eigenvalues of $X$. A similar analysis can be found in [5, Section 10]. By examining the specific structure of the preconditioned matrix, we verify that $X$ satisfies
@@
A detailed proof of this property is presented in Lemma C.1. By [18, Corollary 2.4.4], as long as $2\lVert\Delta E\rVert_{2}<\lambda_{2}$,
@@
By the norm inequalities [22, Table 6.2],
@@
$\begin{split}M_{l}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}\beta^{-1}Y\hat{R}^{-1},\alpha^{1/2}U^{-*}Q\big{\}},\\ M_{r}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{m},\alpha^{-1/2}\beta^{-1}\hat{R}^{-*}Y^{*},\alpha^{1/2}Q^{*}U^{-1}\big{\}},\end{split}$ (5.4)
@@
where $\bar{Z}=[Z,0]\in\mathbb{C}^{m\times n}$. Similarly, the condition number of the preconditioned matrix satisfies
@@
Although it does not necessarily ensure a fast convergence rate theoretically, the condition number of $M_{l}\tilde{F}M_{r}$ has strong implications on the magnitude of the backward and forward errors for the GMRES-based algorithms. For references, a comprehensive bound for the backward and forward errors in the FGMRES algorithm is provided in [9, Section 2]. In practice, the block-diagonal split two-sided preconditioner, due to its smaller condition number and reduced communication cost, shows superior performance in numerical experiments, and therefore emerges as a popular choice for GMRES-based iterative refinement.